OpenStudy (anonymous):

Use the definition of a definite integral (using right end points) to evaluate the integral.. Ive gotten to a certain point and im kinda confused.. (ill type what i have so far below!!).

OpenStudy (anonymous):

\[\int\limits_{0}^{2}(x^2+x)dx =\lim_{n \rightarrow \infty} \sum_{i = 1}^{n} f \frac{2}{n}(\frac{2i}{n})\]\[\lim_{n \rightarrow \infty} \frac{2}{n} \sum_{i=1}^{n}[(\frac{2i}{n})^{2} + \frac{2i}{n}]\] \[\lim_{n \rightarrow \infty}\frac{2}{n} \sum_{i=1}^{n}[\frac{4}{n^2})i^2 + \frac{2}{n}i]\] \[\lim_{n \rightarrow \infty}[\frac{8}{n^3} \sum_{i=1}^{n}i^2 + \frac{4}{n^2} \sum_{i=1}^{n}i]\]

OpenStudy (anonymous):

Any ideas?? :/

OpenStudy (anonymous):

o my god... what math is this?

OpenStudy (anonymous):

Calc 1

OpenStudy (anonymous):

im taking pre cal next semester... i hope its not this complex

OpenStudy (anonymous):

Pre-Calc is Trig and its not bad. Calc 1 isnt really that bad but im just a bit confused lol.

OpenStudy (anonymous):

Wish I could help. I'm really slow when it comes to math.

OpenStudy (anonymous):

Its all good.

OpenStudy (anonymous):

Would appreciate it if you could help me with my prob though

OpenStudy (anonymous):

Ill take a look.